OpenStudy (anonymous):

solve by the elimination method. 9a + 5b = 12 -9a+ b = 24

hartnn (hartnn):

add the 2 equations, what u get ?

hartnn (hartnn):

0+6b= 36 b=...?

OpenStudy (anonymous):

oh nevermind. I just had -6 instead of 6 because I was getting confused when I had to plug that number back in. Thanks

hartnn (hartnn):

but did u get b=6.. ?

hartnn (hartnn):

and a=..?

OpenStudy (anonymous):

yes and then I have 9a+6=12 to find out a

hartnn (hartnn):

* 9a+5*6 =12

hartnn (hartnn):

there is a 5 in 1st equation...

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

so (6,2)

hartnn (hartnn):

i get a=-2

OpenStudy (anonymous):

Thanks so much, I have always struggled with math :/

hartnn (hartnn):

its (6,-2)

OpenStudy (anonymous):

Yes sorry I meant to add - sign

hartnn (hartnn):

then , you are correct.

hartnn (hartnn):

also, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

Thanks again!

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]